# Beyond the Trade-off Curve: Multivariate and Advanced Risk-Utility Maps for Evaluating Anonymized and Synthetic Data

### v2

@@ 2.2.1 Heatmaps
Heatmaps represent the values of a data matrix by encoding them as colour intensities [61]. In the context of statistical disclosure control, they provide a compact overview of how multiple anonymization strategies perform across a range of risk and utility measures.
@@
Two main variants can be distinguished. In the first form, used in Section 3 (cf. Figure 1), the heatmap displays anonymization approaches as rows and evaluation measures as columns. Each tile then shows the performance of one method on one risk or utility metric, with darker or lighter colours indicating higher or lower values, depending on the scale. This layout facilitates visual comparison both across metrics (horizontally) and across anonymization strategies (vertically). When multiple datasets are evaluated, faceting and small multiples can be used to assign one panel per dataset, supporting cross-dataset comparisons. Hierarchical clustering may be applied to reorder rows or columns based on similarity, thereby highlighting patterns or grouping structures.
@@
In an alternative representation, each heatmap corresponds to a single anonymization method, with utility measures along one axis and risk measures along the other. Here, the tile colour reflects the performance on a particular utility–risk pair. This variant highlights joint behaviour across metrics and can be useful for examining interaction structures or identifying imbalanced profiles (e.g., methods that perform well in utility but poorly in specific risk dimensions). However, comparisons between methods are then made across different plots, rather than within a single unified display.
@@
2.2.2 Dot Plots
[…]
Dot plots provide greater accuracy for magnitude comparison, as position encodings are perceptually more effective than colour [2]. They make it easier to spot outliers, skewed profiles, and the relative influence of individual metrics. Pareto-optimal methods can be highlighted using colour or shape coding, and overlays such as risk caps or baseline comparisons can be added for richer interpretation.
2.2.3 Composite Risk/Utility scatterplots
Composite scatterplots summarizes the performance of different anonymization approaches by reducing multiple risk and utility measures into two composite scores [[, see, e.g.,]]Little_2022_Comparing. The horizontal axis shows composite utility e.g. an average of all utility measures, while the vertical axis shows composite risk, similarly an average of all risk measures. Each point in the scatterplot corresponds to an single anonymization approach. The goal of the plot is to highlight trade-offs between disclosure risk and data utility, making it possible to identify anonymization strategies that strike a good balance. The quadrant structure is the same as in a classical R-U map and provides an intuitive interpretation (see Figure 3):
@@ 2.2.3 Composite Risk/Utility scatterplots
The strengths of this visualization are its simplicity and interpretability: it reduces a $n$-dimensional evaluation to two axes, enabling a quick diagnostic view and straightforward comparisons across anonymization strategies.
@@
At the same time, several limitations need to be acknowledged. Collapsing all measures into averages inevitably leads to information loss, hides variability across metrics, and implicitly assumes equal weighting of measures. The approach also ignores metric correlations and it does not display uncertainty such as standard errors or confidence intervals. Finally, results may depend on whether measures have been standardized prior to aggregation.
@@
To enhance interpretability, the composite R–U map can be augmented with several features. First, highlighting the composite Pareto-optimal set separates genuinely efficient anonymization solutions from dominated ones, transforming the plot from a summary into a decision aid. Second, local trade-off annotations reveal the marginal "price" of utility: for each anonymization approach, showing the incremental move to the next better one (e.g., labels with $\Delta$ Utility and $\Delta$ Risk) indicates how much additional disclosure risk is incurred per unit of utility gained. Third, error bars showing the standard deviation of risk and utility along their respective axes indicate dispersion across the underlying measures.
@@
While composite scores are commonly used to summarize performance across multiple risk or utility metrics, their interpretability hinges on the degree to which the constituent measures reflect a coherent underlying construct. To assess this, internal consistency metrics such as Cronbach’s $\alpha$ or McDonald’s $\omega$ can be used as heuristic diagnostics [3, 27, 62, 15]. A value of $\alpha\geq 0.70$ is often cited as indicative of acceptable consistency among items within a block [44, 34, p. 230]. A high value then means that the composite score is a reliable summary of the underlying set of risk or utility measures, reflecting shared variation rather than averaging over unrelated or inconsistent metrics. Note that the mentioned threshold is a rule-of-thumb and context-dependent, and $\alpha$ assumes tau-equivalence – i.e., equal contributions of all indicators to the latent construct – which may not hold in practice. McDonald’s $\omega$ is more general, as it allows heterogeneous loadings across items. In the context of multivariate risk–utility evaluation, internal consistency estimates serve only as rough checks for whether averaging across metrics (to form composite scores) is justifiable.
2.2.4 Parallel Coordinate Plot (PCP)
Parallel coordinate plots [22] are a classical technique for visualizing multivariate data. Shown to be effective for many-objective optimization [32], this visualization likewise applies to the utility–risk formulation. In a PCP, each observation is represented as a polyline intersecting a series of parallel vertical axes, each corresponding to a (min-max normalized) variable (cf. Figure 5). This enables the joint visual inspection of multiple attributes, such as disclosure risk scores and utility indicators across different anonymization methods. PCPs can be useful for identifying outliers, indicating trade-offs between risk and utility, or revealing clusters of methods with similar risk or utility profiles.
2.2.5 Radial Profile Plots (Radar / Origami)
Radial profile plots place each metric on a separate radial axis and connect the values for each anonymization approach or dataset to form a polygonal profile similar to PCP but on a circular basis (cf. Figure 6). Although originally proposed for variables with a circular course [26, p. 78], the visualization is now widely used for general multi-criteria comparison.
One constructs the plot by placing the measure names on the radial axes, plotting the measures’s min–max normalized values as points, and connecting them to form a polygon. This polygonal profile provides a comprehensive view of the performance of an anonymization approach. Area-based summaries of the polygon and its radial profile expose trade-offs: a bulge on one utility axis may coincide with a dip elsewhere, highlighting opposing tendencies. Radar charts provide a holistic view of each anonymization approach’s multivariate measure profile, revealing both how measures vary within an approach and how approaches compare across all measures. However, they can become visually cluttered when many approaches or measures are displayed, inherently assume equal weighting across measures, and produce shapes that depend on the ordering of measures around the circle, which may bias interpretation.
The Origami plot [7] addresses some of these limitations by making the polygonal profile order-invariant through auxiliary axes with points equidistant from the center. In addition, the method allows explicit weighting of variables, so that selected measures can be emphasized more strongly in the visual profile. For a clearer view, one may limit comparisons to Pareto-optimal approaches or a ranking-based subset and group the risk and utility axes and distinguish their labels by color to facilitate interpretation (cf. Figure 6).
2.2.6 Multivariate PCA-based R-U Maps
@@
Multivariate PCA-based R-U maps rely on principal component analysis (PCA) [37, 19], which reduces the dimensionality of multivariate data by projecting it into a low-dimensional space while retaining as much variance as possible. We present two approaches for applying PCA to risk-utility evaluation. The joint PCA approach combines all risk and utility measures in a single analysis and visualizes them in a biplot. The blockwise PCA approach applies PCA separately to risk and utility measure blocks, then plots the resulting principal components in a composite scatterplot.

### v3

@@ 3.1 Heatmaps
As an initial approach to visualizing multiple risk and utility metrics simultaneously, we present heatmaps. Heatmaps represent the values of a data matrix by encoding them as color intensities [Wilkinson_2009_History_heatmap]. In the context of statistical disclosure control, they provide a compact overview of how multiple anonymization strategies perform across a range of risk and utility measures.
@@
Two main variants can be distinguished. In the first form, used in Figure 1, the heatmap displays anonymization approaches as rows and evaluation measures as columns. Each tile then shows the performance of one method on one risk or utility metric (mean of 10 synthetic datasets per SDG), with darker or lighter colors indicating higher or lower values, depending on the scale. This layout facilitates visual comparison both across metrics (horizontally) and across anonymization strategies (vertically). When multiple datasets are evaluated, summarizing like in Figure 1 or faceting and small multiples can be used to assign one panel per dataset, supporting cross-dataset comparisons. Hierarchical clustering may be applied to reorder rows or columns based on similarity, thereby highlighting patterns or grouping structures.
@@
In an alternative representation, each heatmap corresponds to a single anonymization method, with utility measures along one axis and risk measures along the other. Here, the tile color reflects the performance on a particular utility–risk pair. This variant highlights joint behavior across metrics and can be useful for examining interaction structures or identifying imbalanced profiles (e.g., methods that perform well in utility but poorly in specific risk dimensions). However, comparisons between methods are then made across different plots, rather than within a single unified display.
@@
Figure 1 presents a clustered heatmap of min–max normalized measures to reveal cross-method patterns at a glance. To indicate Pareto-optimal SDGs we applied a standard two-objective dominance rule (minimizing Risk, maximizing Utility) explained in Section 2.1. That is, for SDGs $i$ and $j$ (i.e., their mean risk and utility scores over the 10 synthetic datasets), we say that $j$ dominates $i$ if $\mathrm{Risk}_{j}\leq\mathrm{Risk}_{i}\quad\text{and}\quad\mathrm{Utility}_{j}\geq\mathrm{Utility}_{i}$, with at least one strict inequality. SDGs that are not dominated by any other form the Pareto set and are indicated with a star in Figure 1. For the row-wise ordering of the SDGs, hierarchical clustering was used. The heatmap shows that MostlyAI, synthpop, SDV_VAE and simPop are the only SDGs that are on mean basis Pareto-optimal.
Figure 1: Performance of synthetic data generators across utility (bluish) and risk measures (reddish), with values min–max normalized to [0,1] (rows = SDGs, columns = measures). Asterisks indicate composite Pareto-optimal SDGs.
3.2 Dot Plots and Distributional Extensions
Dot plots provide a position-based alternative to heatmaps by encoding values as points on a common scale. Faceting separates risk and utility metrics, and Pareto-optimal methods can be highlighted through shape or color coding. Position-based encodings allow more accurate magnitude comparisons than color-based representations, as perceptual judgments of length and alignment are more reliable than those of hue or intensity [Cleveland_1984_Graphical], facilitating the detection of outliers, skewed profiles, and uneven metric contributions.
When each measure is observed repeatedly (e.g., across multiple synthetic datasets or repeated runs), dot plots can be extended with boxplots to summarize within-method dispersion via the median and interquartile range, revealing instability or occasional extreme outcomes. When only a single value per method is available, the visualization reduces to a classical dot plot. Figure 2 shows this distributional extension.
Figure 2: Distributional dot plot with box plots of risk (left, red) and utility (right, blue) measures across synthesizers. Boxplots display the distribution of individual measure values across replications, while the grey diamonds and connecting lines indicate the per-synthesizer median across measures.
3.3 Composite Risk/Utility Scatterplots
Composite scatterplots summarize the performance of different anonymization approaches by reducing multiple risk and utility measures into two composite scores [[, see, e.g.,]]Little_2022_Comparing. The horizontal axis shows composite utility, e.g., an average of all utility measures, while the vertical axis shows composite risk, similarly an average of all risk measures. Each point in the scatterplot corresponds to an anonymization approach. The goal of the plot is to highlight trade-offs between disclosure risk and data utility, making it possible to identify anonymization strategies that strike a good balance. The quadrant structure is the same as in a classical R-U map and provides an intuitive interpretation (see Figure 3):
@@ 3.3 Composite Risk/Utility Scatterplots
The strengths of this visualization are its simplicity and interpretability: it reduces an $n$-dimensional evaluation to two axes, enabling a quick diagnostic view and straightforward comparisons across anonymization strategies. At the same time, several limitations need to be acknowledged. Collapsing all measures into averages inevitably leads to information loss, hides variability across metrics, and implicitly assumes equal weighting of measures. The approach also ignores metric correlations and it does not display uncertainty such as standard errors or confidence intervals. Finally, results may depend on whether measures have been standardized prior to aggregation.
@@
To enhance interpretability, the composite R–U map can be augmented with several features. First, highlighting the composite Pareto-optimal set separates genuinely optimal anonymization solutions from dominated ones, transforming the plot from a summary into a decision aid. Second, local trade-off annotations reveal the marginal "price" of utility: for each anonymization approach, showing the incremental move to the next better one (e.g., labels with $\Delta$ Utility and $\Delta$ Risk) indicates how much additional disclosure risk is incurred per unit of utility gained. Third, error bars showing the standard deviation of risk and utility along their respective axes indicate dispersion across the underlying measures. Fourth, a horizontal risk-tolerance threshold line (e.g., at a maximum acceptable composite risk level $\bar{r}_{\max}$) can in principle be overlaid to partition the plot into an acceptable region (below the line) and an unacceptable region (above). However, we do not recommend this in practice: individual risk measures rarely have universally agreed thresholds, and aggregating them into a composite further obscures interpretation.
@@
While composite scores are commonly used to summarize performance across multiple risk or utility metrics, their interpretability hinges on the degree to which the constituent measures reflect a coherent underlying construct. To assess this, internal consistency metrics such as Cronbach’s $\alpha$ or McDonald’s $\omega$ can be used as heuristic diagnostics [Cronbach_1951_Coefficient, McDonald_1999_Omega, Zinbarg_2005_Cronbachs, Hayes_2020_Use]. A value of $\alpha\geq 0.70$ is often cited as indicative of acceptable consistency among items within a block [Taber_2018_Use, p. 230]; [Nunnally_1994_Psychometric, p. 245]. A high value then means that the composite score is a reliable summary of the underlying set of risk or utility measures, reflecting shared variation rather than averaging over unrelated or inconsistent metrics. Note that the mentioned threshold is a rule-of-thumb and context-dependent, and $\alpha$ assumes tau-equivalence – i.e., equal contributions of all indicators to the latent construct – which may not hold in practice. McDonald’s $\omega$ is more general, as it allows heterogeneous loadings across items. In the context of multivariate risk–utility evaluation, internal consistency estimates serve only as rough checks for whether averaging across metrics (to form composite scores) is justifiable.
Figure 3 shows an exemplary composite R–U map. For 10 iterations per SDG we compute composite scores by averaging the normalized constituent measures separately for risk (lower is better) and utility (higher is better). The resulting scatterplot places each SDG in the R-U plane. We report Cronbach’s $\alpha$ and McDonald’s $\omega$ for the sets of metrics entering each composite; both indicate that the composites capture a coherent underlying construct. Pareto-optimal SDGs are shown in blue and connected to form the empirical Pareto front. We highlight simPop as the knee point (Section 2.1), defined as the point on the front with the largest perpendicular distance from the straight line joining the two extreme Pareto points, which approximates the location of maximum curvature [Thorndike_1953_Who, p. 275]. Further we display the slope values $\Delta R/\Delta U$, which represent the trade-off from one SDG to another on the Pareto front.
Figure 3: Composite risk vs. utility for SDGs. Pareto-optimal methods are in blue; reliability of composites is indicated by $\alpha$ and $\omega$.
Figure 4: Rays from each SDG mean across 10 iterations to the original $(U_{0},R_{0})$ with labels $\Delta R/\Delta U$ and Euclidean distance $L_{2}$.
In Figure 4 we draw, for each mean SDG point $(U_{i},R_{i})$ across 10 iterations, the ray to the original $(U_{0},R_{0})$ and report the Euclidean $L_{2}$ distance and the slope
interpreted (for $\Delta U>0$) as the incremental risk paid per unit of utility gained relative to the original. A lower $\text{slope}_{i}^{(\mathrm{orig})}$ means a cheaper marginal trade-off from the original, but it is not an overall ranking: global desirability depends on the joint $(U_{i},R_{i})$ and the full metric set. A similar picture is revealed when we look at simple Euclidean distance ($L_{2}$) in the R-U plane. Without explicit scaling/weighting, these can be misleading; for example, SDV_GC and simPop can be equally distant from the original even though the latter Pareto-dominates the former in terms of the mean. We therefore use Pareto dominance to identify “best”, and use slopes only to summarize marginal cost.
3.4 Parallel Coordinate Plot (PCP)
Parallel coordinate plots [[, PCP;]]Inselberg_1985_Plane are a classical technique for visualizing multivariate data and are frequently used in many-objective optimization [Nagar_2023_Visualization]. They are therefore well suited to the joint analysis of disclosure risk and utility measures.
In a PCP, each SDG is represented as a polyline intersecting a series of parallel vertical axes, where each axis corresponds to a min–max normalized evaluation metric (cf. Figure 5). The vertical position on each axis reflects the scaled value of the respective measure. By connecting these positions, the plot encodes the full multivariate performance profile of each method without reducing it to a single aggregate score.
Figure 5 displays the risk and utility measures (scaled to $[0,1]$) across the SDGs. In contrast to the distributional dot plot, which visualizes dispersion across iterations, the PCP shows the mean value over ten synthetic data iterations, resulting in a single aggregated performance profile per SDG. To account for their conceptual and directional differences, risk and utility are shown again in separate facets. Pareto-optimal SDGs are highlighted in color, while non-efficient methods are rendered in neutral grey for contextual comparison.
PCPs are particularly useful for revealing overall performance profiles: relatively flat polylines indicate balanced behavior across metrics, whereas pronounced peaks or troughs suggest specialization or weaknesses in specific dimensions. In contrast to scalar summary indicators, this representation preserves the multivariate structure of the evaluation.
When multiple datasets or repeated runs are analyzed, PCPs can quickly become visually dense if all profiles are overlaid. Two common strategies exist to address this: either aggregating results (e.g., by displaying mean profiles, as done in Figure 5) or employing small multiples by faceting. The former reduces clutter through summarization, while the latter preserves full detail across separate but consistently scaled panels.
A known limitation of PCPs is their sensitivity to axis ordering: different sequences of measures along the horizontal axis can alter visual patterns and emphasis. For this reason, we complement PCPs with order-invariant alternatives such as origami plots (cf. Section 3.5).
3.5 Radial Profile Plots (Radar / Origami)
Radial profile plots represent multivariate performance by placing each metric on a separate radial axis and connecting the corresponding values to form a polygonal profile (cf. Figure 6). Similar in spirit to parallel coordinate plots, they provide a compact visual summary of how an anonymization approach performs across multiple risk and utility dimensions simultaneously. Although early forms were developed for cyclic phenomena [Mayr_1877_Gesetzmaessigkeit, p. 78], radial plots are now widely used for general multi-criteria comparison.
Min–max normalized measures are positioned along equally spaced radial axes and connected to form a polygon. The resulting geometry encodes the multivariate performance profile of a method: differences in radial extent across axes reveal strengths and weaknesses on individual metrics, while the overall shape conveys balance or imbalance across dimensions.
A key limitation of classical radar charts is their sensitivity to the ordering of variables around the circle, as different permutations can produce substantially different visual impressions and alter polygonal areas. In classical radar charts, the polygon area between adjacent axes is jointly determined by neighbouring variables, meaning a variable’s visual contribution depends not only on its own value but also on those of its neighbours — making comparisons sensitive to arbitrary ordering choices. The Origami plot [Duan_2023_Origami] addresses this by introducing auxiliary axes arranged such that each variable’s contribution to the total polygon area is independent of its position around the circle. The plot is therefore only semi-order-dependent: while the overall polygon shape still varies with variable ordering, each individual spike retains the same geometry regardless of where it is placed around the star. This stabilizes area-based comparisons across methods. The framework additionally allows explicit weighting of measures, enabling emphasis on selected criteria when required.
In line with the previous visualizations, we use aggregated performance profiles per SDG. For visual comparability, risk measures (where lower values indicate better performance) are inverted solely for the origami visualization, ensuring that larger radial extensions consistently correspond to more desirable outcomes across both risk and utility dimensions. This transformation is purely visual and does not affect quantitative results.
Because radial plots quickly become cluttered when many profiles are overlaid, we restrict the visualization to the Pareto-optimal SDGs and present pairwise comparisons against the knee solution (cf. Section 2.1). This preserves readability while enabling direct inspection of structural differences between leading methods.
Figure 6 shows these bivariate origami comparisons. The polygonal shapes highlight differences in multivariate composition rather than relying solely on aggregate scores.
Polygon areas can be computed as an additional descriptive summary (cf. Table 3). To ensure interpretability, all axes are harmonized so that larger values indicate better performance. The comparatively small area of the original dataset illustrates the inherent trade-off structure of the evaluation: perfect utility does not offset maximal disclosure risk once axes are aligned, as the origami area reflects balanced performance across all individual measures.
Figure 6: Bivariate origami plots comparing the Pareto-optimal synthesizers against the knee solution (simPop). Each axis represents a normalized performance measure scaled to $[0,1]$, with harmonized orientation such that larger radial distance indicates better performance. Polygon shapes summarize multivariate performance across risk and utility measures, revealing strengths and trade-offs between approaches.
Table 3: Polygon areas of the harmonized origami profiles per synthesizer. Areas are computed from mean-normalized risk and utility measures after aligning axis orientation.
SDG Area
simPop 0.79
MostlyAI 0.75
SDV_VAE 0.73
synthpop 0.73
SDV_GC 0.72
GRETEL 0.68
synthcity 0.66
arf 0.56
original 0.51
3.6 Multivariate PCA-based R-U Maps
@@
Multivariate PCA-based R-U maps rely on principal component analysis [[, PCA;]]1901_Pearson, 1933_Hotelling, which reduces the dimensionality of multivariate data by projecting it into a low-dimensional space while retaining as much variance as possible. We present two approaches for applying PCA to risk-utility evaluation. The joint PCA approach combines all risk and utility measures in a single analysis and visualizes them in a biplot. The blockwise PCA approach applies PCA separately to risk and utility measure blocks, then plots the resulting principal components in a composite scatterplot.
3.6.1 Joint PCA R-U Map
Biplots, introduced by [1971_Gabriel], provide a graphical framework to represent both observations and variables in the same plot. By combining the PCA projection with variable loadings, biplots make it possible to explore patterns, relationships, and group structures in complex datasets. In the context of data anonymization, biplots provide an effective visualization for exploring the trade-off between disclosure risk and data utility while also revealing correlations among multiple evaluation metrics within a single, interpretable representation. Although dimensional reduction methods have previously been used to evaluate the utility of anonymized datasets [[, e.g.,]]2025_Bachot, we are not aware of prior work applying biplots to assess multiple risk and utility measures simultaneously.
A biplot based on a joint PCA of risk and utility measures is presented in Figure 7. To construct the biplot, PCA is applied to the z-standardized risk and utility measures (cf. Section 2.2), where the first two principal components define the axes on which observations and variable loadings are jointly visualized. In the biplot depicted in Figure 7, the first two principal components capture approximately 79% of the total variation. Each plotted point represents an iteration of an anonymization approach, with shape indicating the SDG. The plot can be further augmented with a point color (e.g., to highlight Pareto-optimal approaches) and confidence ellipses. In the biplot, approaches that lie close together have similar risk–utility profiles, while isolated points may indicate outliers (e.g., very low utility and/or very low risk). Loading arrows indicate how the measures correlate, highlighting which risk and utility measures align or contrast. For example, in Figure 7, risk measures (red arrows) are strongly aligned with the PC1 axis, where most risk and utility measures are highly positively correlated, reflecting the risk-utility trade-off. In this illustrative example, generators such as MostlyAI and synthpop are associated with high data utility but also with high risk relative to the other generators. The plot additionally reveals groups of SDGs with similar risk-utility profiles as well as outliers like arf with its particularly low utility regarding the measures Wasserstein and CIProx. When outliers are a concern, a robust PCA variant (e.g., ROBPCA; [Hubert_2005_ROBPCA]) is preferable (see Appendix C for details).
Note that the PCA axes are linear combinations of the original measures and may mix risk and utility; interpretation depends on how the measures load onto the principal components, as well as on the standardization applied and the assumption of linearity.
Additionally, principal components are sign-indeterminate: flipping a component and its loadings leaves the general structure of the biplot unchanged. Although one could fix signs in the joint PCA for readability – e.g., enforce $\operatorname{corr}(PC1,U)\!\geq\!0$ and $\operatorname{corr}(PC2,-R)\!\geq\!0$ and apply the same flips to scores and loadings – we keep the native orientations returned by the algorithms throughout. To aid interpretation, figures annotate the directions of higher utility ($U$) and lower risk ($R$) and display loading arrows.
Compare anonymization approaches across multiple datasets:
When multiple datasets are anonymized (e.g., to compare the performance of different SDGs across datasets), each point in the biplot represents a dataset–anonymization method pair. The biplot then separates datasets and positions anonymization approaches accordingly. When iterations of the same SDG are plotted individually, as in Figure 7, the resulting clusters visualize within-method variation across runs; the same principle extends to multiple datasets, where per-dataset centroids and ellipses can summarize between-dataset variation for each SDG. This complements joint PCA, alignment, and blockwise PCA (see following Section 3.6.2) for multi-dataset comparisons.
PC1 alignment with risk and utility composites:
To support interpretation, we investigate the loadings in detail and relate the PC1 scores to the composite measures of risk and utility (cf. Table 4). The loadings show how risk and utility measures contribute to the principal components, but our broader aim is to assess whether risk and utility as constructs are distinguishable, i.e., whether they occupy separate directions in the principal component space. We therefore correlate the PC scores with the composite risk and utility indicators (each constructed as the mean of their respective measures) to quantify how well each component aligns with these two conceptual dimensions.
Formally, let $X\in\mathbb{R}^{n\times p}$ denote the data matrix where rows correspond to individual SDG iterations (10 runs per SDG) and columns to risk–utility measures, $x_{i}\in\mathbb{R}^{p}$ is the column vector for observation $i$, and $\mu\in\mathbb{R}^{p}$ is the column vector of variable means across all $n$ observations. With $k=2$ retained components, the loading matrix is
$P_{p,k}=[\,\mathbf{p}_{1},\dots,\mathbf{p}_{k}\,]\in\mathbb{R}^{p\times k},$
and the score vector for observation $i$ is
$t_{i}=P_{p,k}^{\top}(x_{i}-\mu)\in\mathbb{R}^{k},$
with PC1 score $t_{i1}=\mathbf{p}_{1}^{\top}(x_{i}-\mu)$. Stacking $t_{i}^{\top}$ as rows yields the score matrix $S\in\mathbb{R}^{n\times k}$.
We then individually align the composites with $t_{1}$ (the vector of first-PC scores across observations). Because Pearson correlation is invariant to positive affine transformations, the choice between z-standardized and original composite scores does not affect $\rho$:
$t_{1}=\beta_{0}+\beta_{U}U+\beta_{R}R+\varepsilon.$
The coefficient of determination $R^{2}$ from this regression expresses the total fraction of variance in the component scores that is explained jointly by the two composites. Thus, while $\rho_{t_{1},U}$ and $\rho_{t_{1},R}$ describe individual associations, $R^{2}$ provides a single summary of joint alignment. For interpretation we report three sets of results: